OpenStudy (anonymous):

ABCD is a rectangle , find the length of each diagonal. AC=2(x-3) BD=x+5

OpenStudy (mertsj):

The diagonals of a rectangle are equal. Write and solve the equation that says so.

OpenStudy (dumbcow):

2x -6 = x+5 x=11 AC=BD = 16

OpenStudy (mertsj):

good job